Solve each equation
2(t-4)= 3-(3t+1)

Answer by stanbon(75887)

Solve each equation
2(t-4)= 3-(3t+1)
2t-8 = 3 -3t - 1
5t = 2+8
5t = 10
t = 2
-----------------
Cheers,
Stan H.